OpenStudy (anonymous):

What is the circumference of a circle with radius 5 feet?

OpenStudy (kendricklamar2014):

Formula:\[C = 2 \pi r\]

OpenStudy (kendricklamar2014):

Now, we substitute what we know:\[C = 2 * 3.14 * 5\]

OpenStudy (k_lynn):

The circumference formula is \[2\Pi r\]Your radius is 5 feet. So plug in 5 for r and solve \[2\times \Pi \times 5\]

OpenStudy (kendricklamar2014):

\[C = 2 \pi r = 31.42\]

OpenStudy (k_lynn):

And \(\pi\) is equal to 3.14, by the way.

OpenStudy (kendricklamar2014):

Exactly ^